# MlingConf: A Comprehensive Study of Multilingual Confidence Estimation on Large Language Models

### v3

@@ 4.1 Results and Analysis on LA Tasks
Despite the powerful multilingual capacity of LLMs, discrepancies exist in the quantity of distinct linguistic training corpora available for each language. The results suggest that ARC. is a metric not significantly related to language usage in LLMs, while the strong performance of ECE in English can be attributed to the extensive training corpus or the specific calibration conducted during training in English. As the only middle-resource language, Thai exhibits a notably lower level of reliability compared to the other high-resource languages.
@@
We also report the response accuracy on LA datasets in Figure 3, English consistently performs better on the datasets exclusively GSM8K. Considering consistency check Table 1, the lowest filter rate in GSM8K translation indicates that mathematical reasoning tasks are minimally affected by language bias. As a result, the performance fluctuations across different languages on GSM8K are relatively small. For Chinese exhibit slightly superior mathematical capabilities compared to English on GSM8K on both GPT-3.5 and Llama-3.1 (Accu. 48.09 and 63.96), it is hypothesized that pre-training corpora contain a substantial amount of Chinese mathematical problems.
Additionally, GPT-3.5 and Llama-3.1 exhibit notable superiority on multilingual tasks over Llama-2 and Vicuna-1.5. Therefore, we only present the results on GPT-3.5 and Llama-3.1 in Table 4 and leave the other two in Appendix E.
@@
4.2 Results and Analysis on LS Tasks
We present the results of the confidence estimation performance on five language-specific subsets of the LSQA dataset using Llama-3.1 as presented in Figure 4 and 5. The key findings and analysis for research question c are listed below.
Recalling research question c, in comparison with the LA task in Sec. 4.1 where linguistic dominance is primarily manifested in English, on LSQA, linguistic dominance is determined by the language of the subset. As illustrated in Figure 4, the diagonal values of the ECE and ARC. heatmaps of Prob. are more pronounced, indicating that when using Prob. confidence, linguistic dominance is more apparent compared to p(True) and Verb.. Consequently, we have opted Prob. for subsequent experiments in 5. Additionally, as shown in Figure 5, although prompting in English performs well and stable across different subsets, there is a noticeable improvement in accuracy when prompts are related to each subset’s language.
The linguistic dominance on LSQA can be conjectured to stem from the fact that such data pertaining to the language-specific cultural, geographical, or social contexts are already included in the pre-training corpora of their respective languages, thereby achieving optimal performance when prompting their specific languages. The use of maximum likelihood estimation during LLM pre-training further accounts for Prob. confidence exhibits more prominent linguistic dominance than p(True) and Verb. confidence estimations.
5 Native-Tone Prompting Method
Confidence estimation performance differences caused by linguistic dominance on LA and LS tasks in Sec. 4 motivate us to further propose a research question How can we better leverage linguistic dominance to improve the LLMs’ reliability and accuracy on LA and LS tasks?
Method
In Sec. 4.1, prompting in English generally outperforms others, hence responding in English on LA tasks can be adequately credible and accurate where linguistic dominance is leading in English. Inspired by the linguistic dominance phenomena on each language-specific LSQA dataset in Sec. 4.2, we propose a simple yet effective Native-Tone Prompting (NTP) strategy to achieve better confidence estimation performance and accuracy on the LS tasks. NTP first prompts LLMs to identify the language context of the question, and then uses that language to answer the question, effectively exhibiting a "native tone" that is more familiar in that language-related context. The prompt of NTP is presented in Appendix C.
@@ 5 Native-Tone Prompting Method
Experiments
In Table 5, we present the results of prompting by any single language versus NTP. The experiments demonstrate that our proposed NTP better leverages the inherent linguistic dominance, thereby yielding more reliable and accurate results than any single language prompt, validating the effectiveness of NTP on LS tasks.
Discussion
Linguistic dominance exerts a significant influence on multilingual confidence estimations on both LA and LS scenarios, where the dominance is manifested in English and query-related language respectively. This phenomenon indicates that the multilingual capabilities and reliability of LLMs are still constrained by the imbalanced training corpus among diverse languages. Furthermore, the reliability and accuracy on English, serving as the primary training corpora, have not been adequately generalized to other languages. Even for semantically equivalent queries, the reliability of responses cannot be consistently maintained, highlighting challenges in enhancing multilingual reliability for LLMs.
@@ 6 Related Works
Both confidence and uncertainty estimations indicate the level of certainty or assurance of a response $\boldsymbol{y}$ generated by LLMs given a query $\boldsymbol{x}$ and are occasionally regarded interchangeably Geng et al. (2023). Previous confidence estimation methods can be categorized into five distinct classes, as illustrated in Figure 6. ➀ Probability-based method: Following model calibration on classification tasks (Guo et al., 2017b), Vazhentsev et al. (2023) intermediately quantify sentence uncertainty over token probabilities; ➁ $p(\mathrm{True})$-based method: Kadavath et al. (2022) instructs the LLM to self-evaluate the correctness of the generated answer by directly accessing $p(\mathrm{True})$; However, both ➀ and ➁ require access to token probabilities and thus are limited to white-box LLMs. ➂ Self-verbalized method: Recently, LLMs’ remarkable instruction-following ability provides a view of expressing confidence in words (Lin et al., 2022; Zhou et al., 2023; Tian et al., 2023a; Xiong et al., 2024) especially for black-box LLMs; ➃ Sampling-based method: By sampling multiple responses to a given question, Xiong et al. (2024); Lyu et al. (2024) aggregate all the confidence scores as the indicator. Kuhn et al. (2023) proposes semantic entropy to quantify uncertainty for sequences with shared meanings at the semantic level. ➄ Training-based method: Mielke et al. (2022); Lin et al. (2022); Han et al. (2024); Kadavath et al. (2022) propose to train an external natural language inference (NLI) module or LLM itself to improve confidence estimations.
Current research for LLM confidence estimations is focused on English. We first present a comprehensive study to investigate LLMs’ confidence estimations in languages other than English. Note that sampling- and training-based methods are omitted in this work as they are cost-expensive and time-consuming for multilingual tasks.
@@
Most recent LLMs primarily pre-trained on English corpora have showcased remarkable capabilities Pires et al. (2019); Shi et al. (2023); OpenAI (2023); Workshop et al. (2023); Chowdhery et al. (2022); Zhang et al. (2022); Brown et al. (2020). However, their efficacy in other low-resource languages remains limited, hindering the global proliferation of LLM deployment and provoking a surge in research interest. Recently, many research works have extended various tasks in multilingual domains such as claim fact-checking (Pikuliak et al., 2023) and jailbreak problem (Deng et al., 2024). Prior studies have also underscored the advantages of leveraging pre-trained multilingual models in diverse cross-lingual applications Wang et al. (2023a, b); Bhattacharjee et al. (2023); Qin et al. (2022). Inspired by them, our research systematically delves into an under-explored domain of multilingual confidence estimation.
@@ 7 Conclusion
This study underscores the necessity of advancing multilingual confidence estimation methods for LLMs to ensure their reliability across diverse linguistic contexts. The proposed MlingConf serves as a valuable and noteworthy benchmark to address the gap in multilingual confidence estimation research. Our findings demonstrate the variability of performance of different confidence estimation methods on both LA and LS scenarios, revealing the influence of linguistic dominance on different tasks. This leads to the NTP strategy, improving accuracy and reliability by aligning the response language with the linguistic context of the query for LS tasks. These insights and the introduction of MlingConf datasets pave the way for future research to refine these methods, enhancing the global applicability and trustworthiness of LLMs.
Limitations and Future Work
@@
Extensions on More Low-Resource Languages
The present study is constrained by the substantial cost associated with the API cost using GPT-4 for translation as well as linguistic verification. This multilingual research is restricted to five languages in the first version. This initial phase aims to delve into confidence estimation within multilingual domains. Our future endeavors will involve the expansion of the benchmark dataset, encompassing additional languages and data points to enrich our investigations.
Extensions on Native-Tone Prompting
@@
In this paper, we introduce several self-constructed multilingual datasets derived from the publicly available dataset. The selection of investigated languages in this work depends on whether we can employ appropriate linguistic experts. Most linguistic specialists are M.Phil. or Ph.D. students majoring in linguistics and others are from crowd-sourcing platforms. We meticulously adhered to legal and ethical standards throughout the data collection process, prioritizing privacy and obtaining informed consent. Linguistic experts were furnished with comprehensive details regarding the study’s objectives, data collection methodologies, and associated risks or benefits. They were afforded the opportunity to seek clarifications and voluntarily provide consent before their involvement. All collected data were exclusively utilized for research purposes.
@@ Appendix C Prompt Details
The translation prompt for multilingual dataset construction and semantic equivalence comparison prompt for consistency check in Sec, 2 are presented in 7 and 8 respectively. Standard multilingual Question-Answering prompts are in 9. Multilingual confidence estimations of P(True) and Verb. are presented in Fig. 10 and 11. Notably, the prompts for self-reflected true probability confidence estimation are followed by previous work Kadavath et al. (2022); Kuhn et al. (2023).
@@ Appendix E Appendix Experiments
We present the experimental results on Llama-2 and Vinuca-1.5 in Table 8. Results suggest that confidence estimation abilities are relatively weak on both Llama-2 and Vinuca-1.5 across three methods.
Dataset Conf. en zh ja fr th
@@
Llama-2
TVQA Prob. 51.92 20.56 51.24 34.12 51.92 32.04 49.51 29.33 49.63 48.72
@@
Vicuna-1.5
TVQA Prob. 45.45 35.34 48.43 47.07 51.75 36.63 46.13 35.19 53.17 40.73

### v4

@@ 4.2 Results regarding Different Languages on LA Tasks
Despite the powerful multilingual capacity of LLMs, discrepancies exist in the quantity of distinct linguistic training corpora available for each language. Results in Table 4 suggest that ARC. is a metric not significantly related to language usage in LLMs, while the strong performance of ECE in English can be attributed to the extensive training corpus or calibrations conducted during training in English. As the only middle-resource language, Thai exhibits a notably lower level of reliability compared to the other high-resource languages.
@@
Considering consistency check in Table 2, the lowest filter rate in GSM8K translation indicates that mathematical reasoning tasks are minimally affected by language bias. As a result, accuracy fluctuations across different languages on GSM8K are relatively small. For that Chinese exhibits slightly superior mathematical capabilities compared to English on GSM8K on both GPT-3.5 and Llama-3.1 (Accu. 48.09 and 63.96), it is hypothesized that pre-training corpora contain a substantial amount of Chinese mathematical problems.
@@
5 Experiments on LS Task
For the LS task, we present the confidence estimation results on five language-specific subsets of LSQA in Figure 4 and 5 in Sec. 5.1. Based on the findings, we then propose a Native-Tone Prompting (NTP) strategy to better leverage linguistic dominance to improve the LLMs’ reliability and accuracy on LS task in Sec. 5.2.
5.1 Results of Different Language-specific Subsets on LS Task
Findings:
Applying prompts in query-related language demonstrates linguistic dominance with better multilingual confidence estimations on LS task. In Figure 4, the diagonal values of the ECE and ARC. heatmaps of Prob. are more pronounced, indicating that when using Prob. confidence, linguistic dominance is more apparent compared to p(True) and Verb.. Consequently, we have opted Prob. for subsequent experiments in Sec. 5.2. Additionally, in Figure 5, although prompting in English performs well and stable across different subsets, there is a noticeable improvement in accuracy when prompts are related to each subset’s language. In comparison with the LA tasks in Sec. 4.1 where linguistic dominance is primarily manifested in English, on LSQA, linguistic dominance is determined by specific language of the subset.
Analysis:
The linguistic dominance on LSQA can be conjectured to stem from the fact that such data pertaining to the language-specific cultural, geographical, or social contexts are already included in the pre-training corpora of their respective languages with higher certainty or confidence, thereby achieving optimal performance when prompting in their respective specific languages.
5.2 Results of Native-Tone Prompting (NTP) Strategy on LS Task
Confidence estimation performance differences caused by linguistic dominance phenomena on the LS task motivate us to explore the improving method. Inspired by results in Sec. 5.1 on each language-specific LSQA subset, we propose a simple yet effective Native-Tone Prompting (NTP) strategy to achieve better confidence estimation performance on the LS task. NTP first prompts LLMs to identify the language context of the question, and then uses that query-related language to answer the question, effectively exhibiting a “native tone” that is more familiar in that language-related context. We present the results of prompting by any single language versus NTP on LSQA in Table 5. The prompt of NTP is presented in the Appendix C.
@@ 5.2 Results of Native-Tone Prompting (NTP) Strategy on LS Task
Findings:
Prompting LLMs using the query-related language can enhance the reliability of confidence estimations and accuracy on LS tasks, which provides an insight to improve LLMs’ reliability regarding the prompt language usage. In Table 5, the experiments demonstrate that NTP better leverages the inherent linguistic dominance, thereby yielding more reliable and accurate results than any single language prompt, validating the effectiveness of NTP on the LS task.
Analysis:
Results in Table 5 indicate that the multilingual capabilities and reliability of LLMs are still constrained by the imbalanced training corpus among diverse languages. The reliability and accuracy on English, serving as the primary training corpora, have not been adequately generalized to other languages. Even for semantically equivalent queries in different languages, the reliability of responses cannot be consistently maintained.
6 Discussion
Extended Confidence Estimations
In Appendix E.2, we further investigate three other confidence estimation methods including 1) paraphrasing the questions; 2) sampling multiple responses (Xiong et al., 2024); and 3) introducing Chain-of-Thought (CoT) (Wei et al., 2022) on both LS and LA tasks. As presented in Table 10 and Figure 13 in the Appendix, all findings on extended three confidence estimations are consistent with previous analysis across all languages. The questions after paraphrasing still maintain semantic equivalence without obvious perturbations for all languages, and LLMs are robust in multilingual confidence estimations to different questions with similar meanings. p(True) and Verb. methods outperform sampling-based methods as the high temperature may incur variability in output spaces which undermines the reliability of QA tasks for all languages. LLMs’ CoT ability can be generalized to multilingual domains, thus benefiting multilingual confidence estimations.
Extended Languages
In Appendix E.3, we also extend the investigations on other five languages derived and translated from TriviaQA into Korean, Arabic, German, Indonesian, and Italian as in Sec. 2. As in Table 11 in the Appendix, linguistic dominance is still performed in English than other languages on the LA task. Low-resource languages demonstrate poor performance in ECE. For the LS task, we also develop a small-size LSQA subset for the above five languages in Table 8 to conduct the NTP method. Experiments suggest that NTP can also generalize and improve the reliability and accuracy in such middle- or low-resource languages.
@@ 7 Related Works
Previous confidence estimation methods can be categorized into five classes, as illustrated in Figure 15 and Appendix F. ➀ Probability-based: Vazhentsev et al. (2023) intermediately quantifies sentence uncertainty over token probabilities; ➁ $p(\mathrm{True})$-based: Kadavath et al. (2022) instructs the LLM to self-evaluate the correctness of the generated answer by directly accessing $p(\mathrm{True})$; Both ➀ and ➁ require access to token probabilities and thus are limited to white-box LLMs. ➂ Self-verbalized: LLMs’ remarkable instruction-following ability provides a view of expressing confidence in words (Lin et al., 2022; Zhou et al., 2023; Tian et al., 2023a; Xiong et al., 2024); ➃ Sampling-based: By sampling multiple responses to a given question, Xiong et al. (2024) aggregates all the confidence scores as the indicator. ➄ Training-based: Lin et al. (2022); Kadavath et al. (2022) propose to train an external module to improve confidence estimations.
@@
Most recent LLMs primarily pre-trained on English corpora have showcased remarkable capabilities Pires et al. (2019); Shi et al. (2023); OpenAI (2023). However, their efficacy in other low-resource languages remains limited. Many research works have extended various tasks in multilingual domains such as claim fact-checking (Pikuliak et al., 2023) and jailbreak problem (Deng et al., 2024). Prior studies have also explored diverse cross-lingual applications Wang et al. (2023a, b); Qin et al. (2022).
@@ 8 Conclusion
This study underscores the necessity of advancing multilingual confidence estimation methods for LLMs to ensure their reliability across diverse linguistic contexts. The proposed MlingConf serves as a valuable and noteworthy benchmark to address the gap in multilingual confidence estimation research. Our findings demonstrate the variability of multilingual confidence estimations on both LA and LS scenarios, revealing the influence of linguistic dominance on different tasks. This leads to the NTP strategy, improving accuracy and reliability by aligning the response language with the linguistic context of the query for LS tasks. These insights and the introduction of MlingConf datasets pave the way for future research, enhancing the global applicability and reliability of LLMs.
Limitations
@@
Expensive Costs to Obtain High-Quality Low-Resource Languages
The present study is constrained by the substantial cost associated with the API cost using GPT-4 for translation as well as linguistic verification. This multilingual research is restricted to five languages in the first version. This initial phase aims to delve into confidence estimation within multilingual domains. Our future endeavors will involve the expansion of the benchmark dataset, encompassing additional languages and datasets to enrich our investigations.
Native-Tone Prompting is a Primary Version
@@
In this paper, we introduce several self-constructed multilingual datasets derived from the publicly available dataset. The selection of investigated languages in this work depends on whether we can employ appropriate linguistic experts. Most linguistic specialists are M.Phil. or Ph.D. students majoring in linguistics and others are from crowd-sourcing platforms. We meticulously adhered to legal and ethical standards throughout the data collection process, prioritizing privacy and obtaining informed consent. Linguistic experts were furnished with comprehensive details regarding the study’s objectives, data collection methodologies, and associated risks or benefits. They were afforded the opportunity to seek clarification and voluntarily provide consent before their involvement. All collected data were solely utilized for research purposes.
Acknowledgements
This work was partially supported by Hong Kong RGC GRF No. 14206324, CUHK direct grant No. 4055209, and CUHK Knowledge Transfer Project Fund No. KPF23GWP20.
@@ Appendix A Language Information
Indonesian id Indo-European
German de Indo-European
Arabic ar Afro-Asiatic
Korean ko Koreanic
Italian it Indo-European
@@ Appendix C Prompt Details
The translation prompt for multilingual dataset construction and semantic equivalence comparison prompt for consistency check in Sec, 2 are presented in 6 and 7 respectively. Standard multilingual Question-Answering prompts are in 8. Multilingual confidence estimations of P(True) and Verb. are presented in Fig. 9 and 10. Notably, the prompts for self-reflected true probability confidence estimation are followed by previous work Kadavath et al. (2022); Kuhn et al. (2023).
@@ Appendix E Appendix Experiments
E.1 Experiments on Llama-2 and Vicuna
We present the experimental results on Llama-2-13B-Chat (Llama-2)  [^7] https://huggingface.co/meta-llama/Llama-2-13b-chat (Touvron et al., 2023), and Vicuna-7B-v1.5 (Vicuna-v1.5)  [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5 (Zheng et al., 2023) in Table 9. Results suggest that confidence estimation abilities are relatively weak in both Llama-2 and Vicuna-1.5 across three methods.
E.2 Experiments of Extended Confidence Estimations
E.2.1 Experiments of Multilingual Confidence Estimations with Paraphrasing
Following Xiong et al. (2024), we investigate the prompt sensitivity for multilingual confidence estimation by introducing perturbations in the questions. We utilize GPT-3.5 to paraphrase the questions in different ways to generate different responses. We sample 200 questions from SciQ and prompt GPT-3.5 to paraphrase these questions. We also employ GPT-3.5 to check the semantic equivalence before and after paraphrasing to ensure the meaning is not changed. The AUROC and ECE results are presented in Table 10 and Figure 13. The findings and analysis are in Sec. 4.1.
E.2.2 Experiments of Multilingual Confidence Estimations of Sampling
To make comparisons, we also present the AUROC and ECE results of sampling-based confidence estimation methods on 200 samples from our multilingual SciQ datasets by setting Temperature T=0.8 on GPT-3.5. We cluster the sampled responses in semantic spaces and calculate the consistency score as Xiong et al. (2024) to represent the confidence. As presented in Table 10 and 13, the results demonstrate that our employed p(True) and Verb. methods outperform sampling-based methods as the high temperature may incur variability in output spaces which undermines the reliability of QA tasks.
E.2.3 Experiments of Multilingual Confidence Estimations using CoT
We supply the Chain-of-Thought (CoT) (Wei et al., 2022) for prompt-based confidence estimations of p(True) and Verbalized methods as in Table 10 and Figure 10. We present the AUROC and ECE results of p(True) and Verb. using CoT on 400 samples from SciQ and LSQA on GPT-3.5. Results suggest that CoT can marginally enhance the reliability of prompt-based confidence estimations in various languages.
Lang. ko id it ar de
Prompt in English
Accu. $\uparrow$ 24.39 40.60 34.58 22.64 54.78
ARC. $\uparrow$ 72.40 70.12 75.45 68.22 76.18
ECE $\downarrow$ 33.55 36.78 33.16 46.78 27.14
NTP Method
Accu. $\uparrow$ 28.60 46.54 39.20 27.44 59.65
ARC. $\uparrow$ 74.66 78.52 77.23 70.17 79.60
ECE $\downarrow$ 28.10 32.44 30.50 42.76 23.18
Table 8: Experimental results of overall Accu., ARC., and ECE on the LSQA dataset by prompting using English and NTP method on other five investigated languages.
Conf. en zh ja fr th
@@ E.2.3 Experiments of Multilingual Confidence Estimations using CoT
TVQA on Llama-2
Prob. 51.92 20.56 51.24 34.12 51.92 32.04 49.51 29.33 49.63 48.72
@@
TVQA on Vicuna-1.5
Prob. 45.45 35.34 48.43 47.07 51.75 36.63 46.13 35.19 53.17 40.73
@@
TVQA on GPT-4o
Prob. 74.22 11.38 72.00 22.34 73.45 20.17 76.18 12.43 78.33 27.48
p(True) 78.15 8.34 77.44 18.24 82.73 16.76 77.24 11.82 83.44 25.15
Verb. 79.33 8.63 78.16 16.33 81.56 17.18 79.68 10.37 86.08 26.44
GSM8K on GPT-4o
Prob. 61.50 21.37 67.34 24.31 63.55 26.47 62.45 28.15 60.44 31.25
p(True) 71.49 18.13 75.19 22.97 73.50 22.77 73.11 21.45 70.60 28.11
Verb. 75.28 16.55 74.30 21.48 75.16 21.64 71.65 23.79 67.54 27.43
CSQA on GPT-4o
Prob. 62.76 21.34 59.46 33.62 58.47 31.78 66.14 28.40 59.19 38.66
p(True) 67.40 17.56 64.75 28.97 63.89 21.40 72.18 19.70 68.56 31.65
Verb. 69.14 16.20 66.13 27.33 66.60 23.65 71.79 19.44 64.33 33.98
SciQ on GPT-4o
Prob. 73.20 27.24 76.44 33.37 79.23 39.16 74.81 31.70 78.20 44.60
p(True) 77.54 16.78 78.62 28.07 83.45 28.07 79.13 21.34 81.39 31.40
Verb. 78.18 17.90 79.13 25.16 84.52 25.44 78.63 20.56 79.27 33.22
Table 9: Experimental results of AUROC (ARC.) and ECE of three confidence estimation methods on four LA datasets on Llama-2 and Vicuna-1.5.
Figure 13: Experimental results of AUROC and ECE of three confidence estimation variants of paraphrasing, sampling, and CoT on LSQA for LS task on GPT-3.5.
Figure 14: Experimental results of AUROC and ECE of verbalized confidence estimation on LSQA for LS task on GPT-4o.
E.3 Experiments on Extended Languages
To further validate the observed linguistic dominance in multilingual confidence estimations, we employ five subsets derived and translated from TriviaQA into Korean (ko), Arabic (ar), German (de), Indonesian (id), and Italian (it) as in Sec. 2. The LA experiments are conducted on dataset translated from TriviaQA in all investigated languages in Table 11. We also develop small-size LSQA subsets for such languages and conduct LS experiments in 8.
Conf. en zh ja fr th
SciQ on GPT-3.5
Prob. 69.58 30.04 67.14 36.77 81.44 45.40 74.35 36.98 72.55 51.34
p(True) 72.80 23.86 77.56 31.99 82.44 38.27 72.00 40.13 63.45 40.80
Verb. 71.43 22.18 72.50 36.47 72.95 31.43 74.16 31.97 73.40 42.34
Re-Prob. 67.47 28.16 72.86 33.43 75.69 41.05 71.40 34.88 80.37 48.96
Re-p(True) 74.14 25.14 82.66 32.04 76.96 36.70 71.48 42.13 64.44 42.05
Re-Verb. 73.80 21.96 73.40 35.13 79.49 30.60 66.16 32.65 73.19 40.44
Sampling 67.55 27.40 71.69 37.97 74.07 42.09 67.94 40.04 66.50 48.65
CoT-p(True) 73.65 22.95 80.05 29.90 82.16 37.10 71.92 30.86 65.90 40.19
CoT-Verb. 73.64 20.60 75.73 32.79 74.61 27.50 72.62 31.26 74.96 40.33
Table 10: Experimental results of AUROC and ECE of several confidence estimation variants of paraphrasing the questions, sampling multiple responses, and adding CoT on SciQ for LA task on GPT-3.5.
E.4 Formatting P(True) Method
The output format issue of the two prompt-based confidence estimation methods is not the primary focus of this study, nor have we observed previous works addressing this problem. However, these issues posed significant challenges during our experiments. With strong instruction-following ability, GPT-3.5 typically generates outputs in the correct format. In cases of occasional formatting errors in the outputs, we employed temperature sampling to re-generate the outputs until the correct format was achieved.
The other three LLMs—Llama-3.1-Instruct, Llama-2-Chat, and Vicuna-v1.5—also demonstrated relatively good adherence to instructions for verbalized confidence estimation. If the correct format was not generated on the first attempt, we also employ temperature sampling multiple times to obtain the expected output.
For the P(True) method, however, output format discrepancies were more pronounced. We explored two approaches: rule-based post-processing and few-shot formatting. Initially, we attempted rule-based post-processing, where the ideal output format would directly consist of “true” or “false” following the input. However, in practice, the models often included their own analysis, and in many cases, the generated sequence did not begin with the desired result. To address this, we detected multilingual keywords within the generated sequence as in Fig. 9.
Since some keywords span more than one token in some languages, we first stored the tokenized sequence before decoding, along with the corresponding logits. We then extracted the logits associated with the tokenized keywords with the normalized or the first-token probability. Despite these efforts, this approach still failed to consistently produce the expected outputs.
We subsequently turned to a few-shot approach. To avoid biases from the order or quantity of “true” and “false” examples in the few-shot samples, we set the number of examples to 10, evenly split between “true” and “false” (five each), and randomized their order in every instance. Ultimately, we found that the few-shot approach not only produced more stable output formats but also yielded more reliable AUROC and ECE results. Therefore, we adopted the few-shot method as our final approach.
Additionally, we considered a training-based method, where negative samples would be constructed to train a classifier head specifically designed to output “true” or “false”. However, this approach was prohibitively costly, as it would require training a separate head for each model in every language. Consequently, we decided against pursuing this method.
Conf. en ko it ar de id
TriviaQA on GPT-3.5
Prob. 69.58 30.04 73.21 46.37 73.08 28.60 71.51 46.78 72.48 33.74 77.37 50.12
p(True) 72.80 23.86 63.19 40.66 70.67 35.47 63.24 50.55 78.49 26.16 66.08 49.81
Verb. 71.43 22.18 72.41 34.80 72.19 41.54 76.65 28.68 68.75 47.14 69.65 47.14
Table 11: Experimental results of AUROC and ECE of confidence estimations on other languages on TriviaQA for LA task on GPT-3.5.
Appendix F Uncertainty Estimations
Both confidence and uncertainty estimations indicate the level of assurance of a response generated by LLMs given a query and are occasionally regarded interchangeably Geng et al. (2023). Uncertainty detection is essential for hallucination mitigation on knowledge-based tasks (Xiong et al., 2024; Varshney et al., 2023; Wang et al., 2024b; Vazhentsev et al., 2023; Wang et al., 2024a; Manakul et al., 2023). To alleviate over-confidence and enhance the reliability of LLMs, reliable uncertainty estimation is essential to determine whether a question is known or not to the LLM (Geng et al., 2023). Both Uncertainty and Confidence estimations can indicate the reliability degree of the responses generated by LLMs, and are generally used interchangeably Xiao et al. (2022); Chen and Mueller (2023); Geng et al. (2023). In this part, we investigate several commonly used confidence & uncertainty estimation methods for generative LLMs as mentioned in Sec. 7. Specifically, we denote $\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ as the confidence score associated with the output sequence $\boldsymbol{y}=[y_{1},y_{2},\dots,y_{N}]$ given the input context $\boldsymbol{x}=[x_{1},x_{2},\dots,x_{M}]$. We also illustrate the summarized estimation methods as well as their disadvantages in Fig. 15.
Likelihood-based Methods:
Following model calibration on classification tasks (Guo et al., 2017b), Vazhentsev et al. (2023); Varshney et al. (2023); Wang et al. (2025) intermediately quantify sentence uncertainty over token probabilities. In traditional discriminative models, except likelihood-based methods, confidence estimations also include ensemble-based and Bayesian methods (Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Xue et al., 2022; Wang and Yeung, 2020; Gal et al., 2016; Abdar et al., 2021; Xue et al., 2021), and density-based methods (Lee et al., 2018). However, this likelihood-based method requires access to token probabilities and thus being limited to white-box LLMs. The likelihood-based confidence is estimated by calculating the joint token-level probabilities over $\boldsymbol{y}$ conditioned on $\boldsymbol{x}$. As longer sequences are supposed to have lower joint likelihood probabilities that shrink exponentially with length, the product of conditional token probabilities of the output should be normalized by calculating the geometric mean by the sequence length Murray and Chiang (2018); Malinin and Gales (2021), and the confidence score can be represented as:
Similarly, the arithmetical average of the token probabilities is adopted in Varshney et al. (2023):
$\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})=\frac{1}{N}{\sum^{N}_{i}p(y_{i}|\boldsymbol{y}_{<i},\boldsymbol{x})}$ (6)
Furthermore, a low probability associated with even one generated token may provide more informative evidence of uncertainty Varshney et al. (2023). Hence, the minimum of token probabilities is also employed.
$\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})={\min\left\{p(y_{1}|\boldsymbol{x}),\dots,p(y_{N}|\boldsymbol{y}_{<N},\boldsymbol{x})\right\}}$ (7)
Prompting-based Methods:
Recently, LLMs’ remarkable instruction-following ability (Brown et al., 2020) provides a view of instructing LLMs to self-estimate their confidence level to previous inputs and outputs including expressing uncertainty in words (Lin et al., 2022; Zhou et al., 2023; Tian et al., 2023a; Xiong et al., 2024), or instructing the LLM to self-evaluate its correctness on $p(\mathrm{True})$ (Kadavath et al., 2022). The $P(\mathrm{True})$ confidence score is implemented by simply asking the model itself if its first proposed answer $\boldsymbol{y}$ to the question $\boldsymbol{x}$ is true Kadavath et al. (2022), and then obtaining the probability $p(\mathrm{True})$ assigned by the model, which can implicitly reflect self-reflected certainty as follows.
$\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})=p(\mathrm{True})=p(\boldsymbol{y}\ \mathrm{is}\ \mathrm{True}?|\boldsymbol{x})$ (8)
Another method is to prompt LLMs to linguistically express tokens of confidence scores in verbalized numbers or words Lin et al. (2022); Mielke et al. (2022); Zhou et al. (2023); Tian et al. (2023b); Xiong et al. (2024).
The sampling-based method refers to randomly sampling multiple responses given a fixed input $\boldsymbol{x}$ using beam search or temperature sampling strategies (Manakul et al., 2023; Xiong et al., 2024; Lyu et al., 2024). Various aggregation methods are adopted on sampled responses to calculate the consistency level as the confidence score. Kuhn et al. (2023) proposes semantic entropy to quantify uncertainty for sequences with shared meanings at the semantic level. Moreover, some uncertainty quantification methods are used to calculate the entropy indicating the dispersion level of multiple outputs (Kuhn et al., 2023; Lin et al., 2023).
Training-based Methods:
For training methods, an external evaluator trained on specific datasets is introduced to output a confidence score given an input and an output. The evaluator can be a pre-trained NLI model (Mielke et al., 2022), or a value head connected to the LLM output layer (Lin et al., 2022; Kadavath et al., 2022), or the LLM itself (Han et al., 2024; Xue et al., 2024).
Figure 15: An illustration of several confidence estimation methods as well as their drawbacks. Note that sampling- and training-based methods are omitted in this work as they are cost-expensive and time-consuming for multilingual confidence estimations. All complete multilingual prompts used in this work are presented in Appendix C. In addition, although confidence and uncertainty are always used interchangeably, the former confidence pertains to the model’s certainty regarding a specific generation, while the latter uncertainty denotes the "dispersion" of potential predictions for a given context. In this work, the semantically equivalent inputs in various languages are thoroughly distinct in the token space. Therefore, we utilize confidence estimation in this work, albeit specific uncertainty quantification methodologies are still applicable.